OpenStudy (anonymous):

what is 3/8 x 2/3 draw a model please

OpenStudy (mathmale):

\[\frac{ 3 }{ 8 }*\frac{ 2 }{ 3 }\]

OpenStudy (jdoe0001):

hmmm I even got them a bit mixed.... hmm

OpenStudy (mathmale):

First, cancel out the "3"s. Write "1"s where the "3"s were. Then, reduce the fraction 2/8. Please show your work.

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{8}\times \cfrac{2}{3}\implies \cfrac{3\times 2}{8\times 3}\implies \cfrac{?}{?}\)

OpenStudy (mathmale):

Cupcake: May we have your attention, please?